# NUDT4MSTAR: A New Dataset and Benchmark Towards SAR Target Recognition in the Wild

### v1

@@ Abstract
Synthetic Aperture Radar (SAR) stands as an indispensable sensor for Earth observation, owing to its unique capability for all-day imaging. Nevertheless, in a data-driven era, the scarcity of large-scale datasets poses a significant bottleneck to advancing SAR automatic target recognition (ATR) technology. This paper introduces NUDT4MSTAR, a large-scale SAR dataset for vehicle target recognition in the wild, including 40 target types and a wide array of imaging conditions across 5 different scenes. NUDT4MSTAR represents a significant leap forward in dataset scale, containing over 190,000 images—tenfold the size of its predecessors. To enhance the utility of this dataset, we meticulously annotate each image with detailed target information and imaging conditions. We also provide data in both processed magnitude images and original complex formats. Then, we construct a comprehensive benchmark consisting of 7 experiments with 15 recognition methods focusing on the stable and effective ATR issues. Besides, we conduct transfer learning experiments utilizing various models trained on NUDT4MSTAR and applied to three other target datasets, thereby demonstrating its substantial potential to the broader field of ground objects ATR. Finally, we discuss this dataset’s application value and ATR’s significant challenges. To the best of our knowledge, this work marks the first-ever endeavor to create a large-scale dataset benchmark for fine-grained SAR recognition in the wild, featuring an extensive collection of exhaustively annotated vehicle images. We expect that the open source of NUDT4MSTAR will facilitate the development of SAR ATR and attract a wider community of researchers.
@@
Synthetic aperture radar (SAR), automatic target recognition (ATR), classification, detection, dataset, benchmark
@@ 1 Introduction
Can humans perceive and recognize objects on the Earth beyond the limitations of optical vision using microwave vision [20]? Synthetic aperture radar (SAR) [72, 61], an active imaging sensor, acquires the ground scattering characteristics by emitting the electromagnetic microwave. It has emerged as an indispensable tool for information acquisition in Earth observation and target surveillance [22, 76, 20], as its resolution remains unaffected by the observation distance or lighting conditions. Moreover, it has the advantage of operating effectively in various weather conditions. SAR interpretation has witnessed remarkable progress, propelled by the advancement of high-resolution imaging and artificial intelligence technologies [20, 64]. SAR automatic target recognition (ATR), which autonomously detects and classifies objects of interest (e.g., vehicles, ships, aircraft, and buildings) within imagery, is a critical frontier in advancing the capabilities of SAR interpretation. It has various civil and military applications such as urban management, disaster assessment, emergency rescue, and global survenlience [44, 18, 49, 25, 98]. Since the 1970s, researchers have been dedicated to developing a generalized stable SAR ATR system. However, the recognition of objects from SAR imagery have encountered significant challenges.
Figure 1: Motivation of our NUDT4MSTAR. Subfigure (a) shows the number of articles on SAR ATR in the last decade. As the first SAR target classification dataset, MSTAR has been the mainstream data benchmark. Subfigure (b) illustrates the data diversity of NUDT4MSTAR compared to MSTAR and its potential to become a new data benchmark.
@@
Special characteristics. SAR imagery has special electromagnetic scattering attributions but is devoid of common features such as color and texture typically present in human vision. The scattering properties of targets are highly influenced by a multitude of factors, including the imaging angle, resolution, polarization, background, and the geometry structure and material of targets themselves [67, 39]. Additionally, SAR images suffer quality issues, such as speckle noise, geometric distortion, and image Defocus. These factors make it difficult to interpret SAR data, a recognition system independent of human vision.
@@
Insufficient samples. Acquiring SAR images encompassing many categories under diverse operating conditions presents a significant challenge, and annotating SAR images is considerably more costly and labor-intensive than RGB images. Three decades have elapsed since the advent of the first SAR classification dataset in the 1990s, the Moving and Stationary Target Acquisition and Recognition (MSTAR) [2] in Fig. 2. The largest publicly accessible SAR target classification dataset remains limited to a small set of no more than 30,000 images. Besides, recent target classification datasets (OpenSARShip [30] and FUSAR-Ship [29]) have severe sample imbalance, ranging from a few to thousands of samples per class. A substantial amount of work remains to be undertaken for SAR target datasets.
@@
Nonstandard benchmarks. A standardized dataset and experimental settings serve as the north star for developing advanced SAR ATR techniques. However, the mainstream [^1] For counting SAR ATR papers, we use the search formula “SAR target recognition (Title) or SAR ATR (Title) or SAR target classification (Title) or SAR classification (Title) and radar (Topic) ” on Web of Science. We added “and MSTAR (Topic)” to the previous search formula to count the number related to MSTAR. MSTAR dataset in Fig. 1 has dozens of distinct experimental settings, and existing algorithms have nearly reached near-saturation performance on this ideal benchmark [39]. Other classification datasets, such as OpenSARShip and FUSAR-Ship, also exhibit different experimental settings and dataset splitting. Moreover, the limiting access to non-open-source dataset settings and methods further intensifies this challenge.
@@
Researchers have struggled to construct a large number of SAR target datasets to study generalized stable SAR ATR algorithms. Over the past few years, there has been a remarkable surge in remote sensing datasets across various modalities, including visible light [4, 51], SAR [52, 85], infrared [90], and multimodal [58]. As depicted in Fig. 2, recent SAR classification datasets have significantly improved data diversity. In particular, SAR detection datasets [52, 85] with 100,000 images have emerged. Yet, regarding SAR target recognition, no dataset has emerged that can completely surpass the MSTAR benchmark dataset, released in the 1990s, regarding the diversity of target types, imaging conditions, and data formats it offers. Notably, the lack of open-source availability, inadequate samples, long-tailed distributions, unabundant collection conditions, and independent experimental settings are still all issues that warrant attention. We suggest that a large-scale fine-grained SAR target dataset and benchmarks to investigate the electromagnetic scattering properties of targets across diverse conditions represent a critical and enduring endeavor.
Figure 3: NUDT4MSTAR. We collect target samples from 21 classes and 40 types of targets in 5 scenes with 3 modalities of sensors. Detailed target information is illustrated in Fig. 7. Due to the limitation of the scene size, there are two settings for target deployment, the first with 40 types in city, factory, and sandstone; the second with 11 types in woodland and bare soil. These vehicle types cover most current civilian vehicles in structure, size, and usage. In difficult scenarios (city and factory), vehicle positions are adjusted during acquisition to avoid pseudo-correlation between the target and the background. Although we focus on target characteristics in the SAR sensor, we also collect data from other sensors for future analysis of their differences.
@@
As a first step towards building a large-scale SAR target dataset, we introduce NUDT4MSTAR, a fine-grained dataset for SAR vehicle target recognition in Fig. 3. In particular, NUDT4MSTAR collects 194,324 object images in the wild from 40 target types, 5 scenes, and various imaging conditions with detailed annotation [^2] This dataset focuses on the fine-grained single target recognition problem in SAR ATR, i.e., the image classification task, but we provide a horizontal bounding box label for the detection task.. It is the largest public SAR vehicle recognition dataset, even 10 times the size of any similar vehicle dataset previously available in Fig. 2. Then, to facilitate innovation and comparisons, we build a well-designed benchmark with 7 experimental settings and 15 recognition methods. Besides, the transfer learning results of various models trained on this dataset to other datasets show that our fine-grained recognition task benefits the feature learning of different ground targets. Finally, We provide detailed analyses and discussions to explore important issues for future research.
@@
Our work shows the huge challenges and potential in SAR recognition and provides the benchmark for stable fine-grained ATR algorithms. The dataset and benchmark are available on our GitHub at https://github.com/waterdisappear/NUDT4MSTAR. We believe that our new dataset, with comprehensive imaging conditions and experiment benchmark, will further promote the promise and reproducible research in SAR ATR. The main contributions of this paper are three points:
@@
1. We build the largest dataset for SAR vehicle target recognition in microwave vision. With various target types and imaging conditions, NUDT4MSTAR not only provides a systematic study of the effects of different conditions on SAR target characteristics but also promotes the development of new ATR technologies and research topics.
2. We constructed a target classification system for this vehicle dataset, including 4 categories, 21 classes, and 40 types. A transfer learning experiment based on pre-training on this dataset shows that a large fine-grained task can provide good weight initializations for recognizing ground targets.
3. We build a comprehensive benchmark with 7 experimental settings and 15 recognition methods. This benchmark provides detailed analysis and discussion that can guide SAR recognition algorithms’ designs and facilitate reproducible comparisons.
@@ 2 Related Work
TABLE I: Statistics of the SAR target classification datasets from the 1990s to 2020s. We focus on the electromagnetic scattering properties of individual objects, i.e., the most relevant to the classification dataset and task. Existing SAR target classification datasets face challenges such as no open source, insufficient samples, long-tailed distributions, and single collection conditions. OA.: Open access. # Classes.: Number of object classes. # Types.: Number of object types. Img. Size: image size. # Img.: Number of images. Res.: Resolution. Pol.: Polarization. “-” refers to unknown.
@@
2.1 Datasets for SAR target classification
Datasets are the cornerstone of developing new techniques for SAR ATR, especially in a deep learning-based era. In the past decades, the MSTAR [2] dataset has played a key role in SAR target classification tasks. However, current deep learning-based methods are saturated on this dataset benchmark setting. MSTAR’s insufficient data volume also suffers from data biases such as background correlation [39, 47]. Therefore, many datasets have been built to advance the development of this field, especially in recent years. In order to provide a comprehensive analysis, we discuss SAR target classification datasets that focus on individual target characteristics from the 1990s to the 2020s, which are detailed listed in Table I.
@@ 1. $\bullet$
    Richness. The dataset should include rich target categories, scenes, and imaging conditions to approximate the real world and reduce data bias.
@@ 2. $\bullet$
    Metadata. The dataset needs enough metadata, such as complex data, scene description, and imaging parameters. Researchers can propose new algorithms by combining metadata with SAR image characteristics.
  3. $\bullet$
    Standardization. An official benchmark and setting are beneficial to fair comparisons across different methods. In addition, the dataset format should be compatible with mainstream datasets for expansion and scaling.
@@ 2.2 Review of SAR Target Classification Algorithms
Traditional features. Previous studies have proposed a plethora of valuable features, encompassing geometric structure features [33, 23], electromagnetic-scattering features [63, 37, 11, 93, 45], time-frequency features [71], local descriptors [13, 70], and sparse representations [28]. These features are particularly tailored to the special coherent imaging mechanisms. For instance, SAR-HOG [70] employs a ratio-based gradient computation rather than the traditional differential gradient approach, and Attributed Scattering Centers (ASC) provides a feature set including location, geometry, and polarization information. Notably, extracting electromagnetic-scattering features relies on complex images with phase information, underscoring the necessity of providing comprehensive data formats.
Deep learning. Deep learning algorithms excel at uncovering intricate patterns within datasets, particularly when ample data is available. Nowadays, these algorithms achieve a remarkable nearly 99% accuracy on MSTAR’s basic Standard Operating Condition (SOC) settings, but the Extended Operating Condition (EOC) with robustness and few-shot are more challenging problems. Convolutional Neural Network (CNN) [40, 57] is a common network in SAR target classification. All-convolutional network (A-ConvNet) [8] was proposed to reduce overfitting and had high accuracy under many experiment settings but is not robust to noise. In recent years, Transformer [78, 14] has also been used in SAR target classification. After a decade of significant progress, SAR target recognition now incorporates many advanced deep learning techniques such as data augmentation [12, 41, 24], attention mechanism [65, 80], domain alignment [50], reinforcement Learning [35, 7], transfer learning [89], meta-learning [21], physical deep learning [9, 82, 32, 53], semi-supervised learning [97], and self-supervised learning [92, 34, 48]. The majority of research endeavors are based on the MSTAR dataset, and a subset of studies also create the different non-open source settings of OpenSARShip and FUSAR-Ship datasets. Nonetheless, the reliance on outdated datasets and the lack of standardized benchmarks are impeding the advancement of SAR target recognition in the 2020s.
@@
3 NUDT4MSTAR
TABLE II: Statistics of NUDT4MSTAR. We segment the targets with fixed-size slices and random offsets to investigate their characteristics and aim for robust recognition under different collection conditions. We varied the target location in urban and factory scenes and collected data at the factory several times. In addition, some targets can not be labeled due to heavy occlusion. # Types.: Number of object types. Res.: Resolution. Pol.: Polarization. Dep.: Depression angle. Azi.: Target azimuth angle interval. Img. Size: image size. # Img.: Number of images in ground and slant range coordinate systems.

### v2

# NUDT4MSTAR: A Large Dataset and Benchmark Towards Remote Sensing Object Recognition in the Wild

@@ Abstract
As an indispensable sensor for Remote sensing, Synthetic Aperture Radar (SAR) has a unique capability for all-day imaging. Nevertheless, in a data-driven era, the scarcity of large-scale datasets poses a significant bottleneck to advancing SAR automatic target recognition (ATR) technology. This paper introduces NUDT4MSTAR, a large-scale SAR dataset for remote sensing target recognition in the wild, including 40 vehicle target types and various imaging conditions across 5 realistic scenes. NUDT4MSTAR represents a significant leap forward in dataset scale, containing over 190,000 images—tenfold the size of its predecessors. We meticulously annotate each image with detailed target information and imaging conditions. Besides, data in both processed magnitude images and original complex formats are provided. Then, we construct a comprehensive benchmark consisting of 7 experiments with 15 recognition methods focusing on the stable and effective ATR issues. Besides, we conduct transfer learning experiments utilizing various models training on NUDT4MSTAR and apply them to three other target datasets, demonstrating its substantial potential for the broader field of ground objects ATR. Finally, we discuss this dataset’s application value and ATR’s significant challenges. To the best of our knowledge, this work marks the first-ever endeavor to create a large-scale dataset benchmark for fine-grained SAR recognition in the wild, featuring an extensive collection of exhaustively annotated vehicle images. We expect that the open source of NUDT4MSTAR will facilitate the development of SAR ATR and attract a wider community of researchers.
@@
Remote sensing, transformer, benchmark dataset, performance evaluation, synthetic aperture radar, automatic target recognition, image classification, object detection, foundation model, deep learning
@@ 1 Introduction
Figure 2: Motivation of our NUDT4MSTAR. Subfigure (a) depicts the most frequent keywords in 21,780 journal papers published in the remote sensing field (TGRS, J-STARS, GRSL, ISPRS P&R, and JAG) from 2020 to 2024. The size of each word is proportional to its frequency, highlighting that concepts such as synthetic aperture radar (SAR), image classification, and object detection [21, 10, 99] have garnered substantial attention. Subfigure (b) focuses on the number of publications related to SAR ATR over the past five years, a cross-area of the concepts highlighted in Subfigure (a). As the pioneering dataset for SAR target classification, MSTAR has long served as the predominant benchmark due to its unique data diversity and accumulated benchmarks. However, the lack of large-scale datasets has significantly limited the growth of this research field in recent years.
Remote sensing imagery interpretation is crucial for understanding our Earth comprehensively via diverse multi-modal sensors [64, 27, 111]. Recently, the advent of the big data era has propelled the evolution of pre-trained foundation models, which have been widely applied across various remote sensing downstream tasks involving different multi-modal data [30, 107, 87, 26, 3, 58]. Nevertheless, not all modal tasks in remote sensing possess a large number of target samples to exploit the potential of big data and foundation models.
@@
Synthetic aperture radar (SAR) [78, 36], an active imaging sensor, acquires the ground scattering characteristics by emitting the electromagnetic microwave. As shown in Fig 2 (a), it has emerged as an indispensable tool for information acquisition in Remote sensing and Earth observation [23, 83, 21], as its resolution remains unaffected by the observation distance or lighting conditions. SAR automatic target recognition (ATR), which autonomously detects and classifies objects of interest (e.g., , vehicles, ships, aircraft, and buildings) within imagery, is a critical increasing frontier in SAR interpretation, as illustrated in Fig. 2 (b). It has various civil and military applications, such as urban management, disaster assessment, emergency rescue, and global surveillance [10, 54, 28, 108]. Can humans perceive and recognize objects on the Earth beyond the limitations of optical vision using microwave vision [21, 36, 49]? Since the 1970s, researchers have been dedicated to developing a generalized stable SAR ATR system. However, recognizing objects within SAR images has encountered significant challenges in a big data era, especially the scarcity of large-scale target samples.
@@
Special characteristics. SAR imagery has special electromagnetic scattering attributions but is devoid of common features such as color and texture typically present in human vision. The scattering properties of targets are highly influenced by a multitude of factors, including the imaging angle, resolution, polarization, background, and the geometry structure and material of targets themselves [73, 44, 69]. Additionally, SAR images suffer quality issues, such as speckle noise, geometric distortion, and image Defocus. These factors make it difficult to recognize objects in SAR, a unique recognition system independent of human vision.
@@
Insufficient samples. Acquiring SAR images encompassing many categories under diverse operating conditions presents a significant challenge, and annotating SAR images is considerably more costly and labor-intensive than RGB images. Three decades have elapsed since the advent of the first SAR classification dataset in the 1990s, the Moving and Stationary Target Acquisition and Recognition (MSTAR) [2]. As shown in Fig. 3, the largest publicly accessible SAR target classification dataset remains limited to a small set of fewer than 30,000 images. Besides, recent classification datasets (OpenSARShip [34] and FUSAR-Ship [33]) have severe sample imbalance, ranging from a few to thousands of samples per class. A substantial amount of work remains to be undertaken for SAR target datasets.
@@
Nonstandard benchmarks. A standardized dataset and experimental settings serve as the north star for developing advanced SAR ATR techniques. However, the mainstream MSTAR dataset in Fig. 2 (b) has dozens of distinct experimental settings, and existing algorithms have nearly reached near-saturation performance on this ideal benchmark [44]. Other classification datasets, such as OpenSARShip and FUSAR-Ship, also exhibit different experimental settings and dataset splitting. Moreover, the limiting access to non-open-source dataset settings and methods further intensifies this challenge.
@@
Researchers have struggled to construct many SAR target datasets to study generalized stable SAR ATR algorithms. Over the past few years, there has been a remarkable surge in remote sensing datasets across various modalities, including visible light [5, 56], SAR [57, 93], infrared [98], and multimodal [65]. Recent SAR classification datasets in Table II have significantly improved data diversity. In particular, SAR detection datasets [57, 93] with 100,000 images have emerged. Yet, regarding SAR target recognition, no dataset has emerged that can completely surpass the MSTAR benchmark dataset, released in the 1990s, regarding the diversity of target types, imaging conditions, and data formats it offers. Notably, the lack of open-source availability, inadequate samples, long-tailed distributions, unabundant collection conditions, and independent experimental settings are still all issues that warrant attention. We suggest that a large-scale fine-grained SAR target dataset and benchmarks to investigate the electromagnetic scattering properties of targets across diverse conditions represent a critical and enduring endeavor.
TABLE I: MSTAR v.s. NUDT4MSTAR. We can see the potential of the NUDT4MSTAR dataset as a new benchmark, surpassing MSTAR in terms of data diversity and challenges. AFRL: Air Force Research Laboratory. NUDT: National University of Defense Technology. # Types.: Number of object types. # Samples: Number of samples.
Dataset MSTAR NUDT4MSTAR
Year 1995 2025
Country USA CHN
Institution AFRL NUDT
# Types 10 40
Location ideally centered random, noncentered
Scene grass 5 realistic scenes
Platform Airborne Airborne
[…]
Band X X, Ku
Polarization single quad
Depression angle () 15, 17, 30, 45 15, 30, 45, 60
[…]
Annotation classification classification, detection
# Samples 14,557 194,324
@@
As a first step towards building a large-scale SAR target dataset, we introduce NUDT4MSTAR [^1] The name NUDT4MSTAR comes from the abbreviation of the College of Electronic Science and Technology, the National University of Defense Technology, as Department IV. This dataset focuses on the finegrained single target recognition problem in SAR ATR, i.e., the image classification task, but we provide a horizontal bounding box label for the object detection task. NUDT4MSTAR: https://github.com/waterdisappear/NUDT4MSTAR., a fine-grained dataset for vehicle target recognition as a new breakthrough to the previous benchmark dataset MSTAR in Table. I. In particular, NUDT4MSTAR collects 194,324 object images in the wild from 40 target types, 5 scenes, and various imaging conditions with detailed annotation. It is the largest public SAR vehicle recognition dataset, even 10 times the size of any similar vehicle dataset previously available, as shown in Table. II. Then, to facilitate innovation and comparisons, we build a well-designed benchmark with 7 experimental settings and 15 recognition methods. Besides, the transfer learning results of various models trained on this dataset to other datasets show that our fine-grained recognition task benefits the feature learning of different ground targets. Finally, We provide detailed analyses and discussions to explore important issues for future research.
@@
Our work shows the huge challenges and potential in SAR recognition and provides the benchmark for stable fine-grained ATR algorithms. The dataset and benchmark are available on our GitHub. We believe that our new dataset, with comprehensive imaging conditions and experiment benchmark, will further promote the promise and reproducible research in SAR ATR. The main contributions of this paper are three points:
@@
1. We build a large dataset for remote sensing object recognition in the Wild. NUDT4MSTAR includes 40 distinct vehicle target types and over 194,324 target samples across 5 realistic scenes and various imaging conditions. It provides a systematic study of the effects of different conditions on SAR target characteristics and promotes the development of new ATR research topics.
2. We provide a comprehensive benchmark with 7 experimental settings and 15 recognition methods, covering both classification and detection recognition tasks. This benchmark offers a standardized experimental platform for researchers to compare the performance of different algorithms, facilitating innovation and development in SAR ATR technology.
3. We demonstrate the NUDT4MSTAR dataset’s potential for foundation models. A transfer learning experiment based on various models pre-trained on this dataset shows that it is not only suitable for vehicle target recognition but also has the potential to support research in other related objects, promoting the application of SAR ATR in multiple situations.
@@ 2 Related Work
2.1 Datasets for SAR target classification
Datasets are the cornerstone of developing new techniques for SAR ATR, especially in a deep learning-based era. In the past decades, the MSTAR [2] dataset has played a key role in SAR target classification tasks. However, current deep learning-based methods are saturated on this dataset benchmark setting. MSTAR’s insufficient data volume also suffers from data biases such as background correlation [44, 52]. Therefore, many datasets have been built to advance the development of this field, especially in recent years. In order to provide a comprehensive analysis, we discuss SAR target classification datasets that focus on individual target characteristics from the 1990s to the 2020s, which are detailed listed in Table II and Fig. 3.
TABLE II: Statistics of the SAR target classification datasets from the 1990s to 2020s. We focus on the electromagnetic scattering properties of individual objects, i.e., the most relevant to the classification dataset and task. Existing SAR target classification datasets face challenges such as no open source, insufficient samples, long-tailed distributions, and single collection conditions. OA.: Open access. # Classes.: Number of object classes. # Types.: Number of object types. Img. Size: image size. # Img.: Number of images. Res.: Resolution. Pol.: Polarization. “-” refers to unknown.
@@ 1. $\bullet$
    Richness. The dataset should include rich target categories, scenes, and imaging conditions to approximate the real world and reduce data bias. Besides, the dataset needs enough metadata, such as complex data, scene description, and imaging parameters. Researchers can propose new algorithms by combining metadata with SAR image characteristics.
@@ 2. $\bullet$
    Standardization. An official dataset and benchmark are beneficial to fair research and comparisons. In addition, the dataset format and target system should be standardized to be compatible with mainstream datasets and for further expansion and scaling.
@@ 2.2 Review of SAR Target Classification Algorithms
Traditional features. Previous studies have proposed a plethora of valuable features, encompassing geometric structure features [38, 24, 13], electromagnetic-scattering features [70, 42, 102, 50], time-frequency features [77], local descriptors [15, 76], and sparse representations [32]. These features are particularly tailored to the special coherent imaging mechanisms. For instance, SAR-HOG [76] employs a ratio-based gradient computation rather than the traditional differential gradient approach, and Attributed Scattering Centers (ASC) provides a feature set including location, geometry, and polarization information. Notably, extracting electromagnetic-scattering features relies on complex images with phase information, underscoring the necessity of providing comprehensive data formats.
Deep learning. Deep learning algorithms excel at uncovering intricate patterns within datasets, particularly when ample data is available. Nowadays, these algorithms achieve a remarkable nearly 99% accuracy on MSTAR’s basic Standard Operating Condition (SOC) settings, but the Extended Operating Condition (EOC) with robustness and few-shot are more challenging problems. Convolutional Neural Network (CNN) [45, 63] is a common network in SAR target classification. All-convolutional network (A-ConvNet) [9] was proposed to reduce overfitting and had high accuracy under many experiment settings but is not robust to noise. In recent years, Transformer [85, 16] has also been used in SAR target classification. After a decade of significant progress, SAR target recognition now incorporates many advanced deep learning techniques such as data augmentation [14, 46, 25], attention mechanism [71, 88], domain alignment [55], reinforcement Learning [40, 8], transfer learning [97], meta-learning [22], physical deep learning [11, 90, 37, 59], semi-supervised learning [106, 80], and self-supervised learning [101, 39, 53]. The majority of research endeavors are based on the MSTAR dataset, and a subset of studies also create the different non-open source settings of OpenSARShip and FUSAR-Ship datasets. Nonetheless, the reliance on outdated datasets and the lack of standardized benchmarks are impeding the advancement of SAR target recognition in the 2020s.
@@
3 Building NUDT4MSTAR: Why, What, How
In this section, we present NUDT4MSTAR from three perspectives: (1) Why we constructed this dataset in the 2020s; (2) What we did to ensure its diversity and standardization; (3) How the diversity and statistical analysis of this dataset to support SAR ATR.
TABLE III: Statistics of NUDT4MSTAR. We segment the targets with fixed-size slices and random offsets to investigate their characteristics and aim for robust recognition under different collection conditions. We varied the target location in urban and factory scenes and collected data at the factory several times. In addition, some targets can not be labeled due to heavy occlusion. # Types.: Number of object types. Res.: Resolution. Pol.: Polarization. Dep.: Depression angle. Azi.: Target azimuth angle interval. Img. Size: image size. # Img.: Number of images in ground and slant range coordinate systems.